a market researcher believes that industry type is independent of the operating margin. He compiles a table with seven industry classifications and classifies operating margin into five levels. When applying a chi-square test of independence to this table, the number of degrees of freedom is
Answer #1

The number of degrees of freedom = (r - 1)*(c - 1)

r = 7

c = 5

The number of degrees of freedom = (7 - 1)*(5 - 1) = 6*4 = 24

Please give me a thumbs-up if this helps you out. Thank you!